OpenStudy (anonymous):

Help please: ================ Solve the system of equation for positive real numbers: (1/xy) = (x/z) + 1,(1/yz)=(y/x)+1,(1/zx)=z/y+1

OpenStudy (raden):

assume x=y=z, with x,y, and z are real positive numbers for the 1st equation, (1/xy) = (x/z) + 1 can rewrite be : 1/(x*x) = (x/x) + 1 or 1/x^2 = 1+1 1/x^2 = 2 x^2 = 1/2 x=+- (sqrt(1/2)) (but - sign not satisfied) so, x=sqrt(1/2) because x=y=z, obviously y=sqrt(1/2) , and z=sqrt(1/2) also

OpenStudy (anonymous):

Thanks RadEn,it is working.However , how we can justify the assumption of x=y=z?